OpenStudy (anonymous):

Choose a variable to isolate and what it is equal to in the system of equations below. 3x=y+8 2y=4x-3 a.y=3x+8 b.2y=4x-3 c.y=3x-8 d.3y=x-8

OpenStudy (anonymous):

First step: solve either 3x=y+8 or 2y=4x-3 for one variable Which one would you like to solve for?

OpenStudy (anonymous):

2y=4x-3 please

OpenStudy (anonymous):

Ok, do you want to try solving that for y?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

K :) let me know what you get

OpenStudy (anonymous):

i got 6 and 8. i think its wrong

OpenStudy (anonymous):

Ok, you're solving 2y=4x-3 for y, right? We won't be able to get a numerical answer because there are two unknowns here, so we can only solve for it in terms of x. Right now y is being multiplied by 2, so in order to undo the multiplication, we have to divide both sides of the equation by 2. y=4/2x-3/2 or y=2x-3/2

OpenStudy (anonymous):

i think its y=2x-3/2 if you divide 3 by 2 you get 1.5

OpenStudy (anonymous):

Right. I'm a little confused by the question, is it for online math homework? I just noticed that one of your options (b) is the same as one of the equations that they give you.

OpenStudy (anonymous):

yes its online homework. thanks friend =)

OpenStudy (anonymous):

My pleasure, do you think you can get the final answer from here?

OpenStudy (anonymous):

umm idk but im gonna go for b:)

OpenStudy (anonymous):

b might work because it's identical to one of the original equations. There's another choice that might work too. If you solve the top equation (3x=y+8) for y, you get y=3x-8, which is (c)

OpenStudy (anonymous):

Thanks Again Friend =)

OpenStudy (anonymous):

You're welcome, good luck :)